Circular No. 7383

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS Further to IAUC 7376, D. A. Biesecker reports observations (measures

by D. Hammer and himself, reductions by B. G. Marsden) of four more tailless, Kreutz sungrazing comets seen in SOHO C3 data, discovered by K. Schenk, except for C/1999 R4, which was first noted by T. Lovejoy via the SOHO Web page; astrometry and orbital elements appear on the MPECs indicated below.

1999 UT α_{2000} δ_{2000} MPEC Comet
Sept. 5.846 10^{h}36.^{m}0 + 5^{o}48′ 2000-F09  C/1999 R4
20.638 11 29.6 + 0 13  2000-F10  C/1999 S5
21.179 11 34.6 + 0 20  2000-F11  C/1999 S6
24.196 11 42.6 − 1 12  2000-F12  C/1999 S7

COMET C/1999 S4 (LINEAR) The ephemeris below was computed from orbital elements on MPC

38308. As this comet may be making its first pass through the inner solar system (original 1/a = −0.000004 ± 0.000007 AU^{−1}), it should be noted that the observed peak magnitude in July may well be a couple of magnitudes fainter than that given below (which assumes brightening as r^{−4}).

2000TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 16 1^{h}57.^{m}50 +29^{o}43.^{′}8 2.366 1.500 24.^{o}0 15.^{o}9 10.6
21 2 00.04 +30 25.8 2.253 1.430 27.3  18.9 10.3
26 2 02.66 +31 13.1 2.131 1.361 30.7  22.3 10.0
31 2 05.37 +32 07.0 1.999 1.293 34.2  26.2  9.6
June 5 2 08.25 +33 09.1 1.858 1.225 37.7  30.5  9.2
10 2 11.40 +34 21.8 1.708 1.158 41.3  35.3  8.8
15 2 15.01 +35 48.8 1.550 1.093 44.7  40.8  8.3
20 2 19.46 +37 35.7 1.383 1.031 48.0  47.1  7.8
25 2 25.44 +39 52.0 1.209 0.972 50.9  54.3  7.3
30 2 34.40 +42 53.2 1.028 0.917 53.3  62.7  6.7
July 5 2 49.67 +47 06.6 0.846 0.869 54.7  72.7  6.0
10 3 20.64 +53 17.0 0.666 0.828 54.2  85.1  5.3
15 4 41.49 +61 44.0 0.503 0.796 50.3 100.5  4.5
20 8 16.04 +63 08.1 0.392 0.775 42.6 117.4  3.9
25 11 01.02 +41 05.5 0.385 0.765 40.6 120.3  3.8
2000 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green